OpenStudy (anonymous):

Find the solutions x^2 + 2x = 4

OpenStudy (anonymous):

get the equation = to 0

OpenStudy (anonymous):

after that foil it and your done

OpenStudy (anonymous):

Simplifying x2 + 2x = 4 Reorder the terms: 2x + x2 = 4 Solving 2x + x2 = 4 Solving for variable 'x'. Reorder the terms: -4 + 2x + x2 = 4 + -4 Combine like terms: 4 + -4 = 0 -4 + 2x + x2 = 0 Begin completing the square. Move the constant term to the right: Add '4' to each side of the equation. -4 + 2x + 4 + x2 = 0 + 4 Reorder the terms: -4 + 4 + 2x + x2 = 0 + 4 Combine like terms: -4 + 4 = 0 0 + 2x + x2 = 0 + 4 2x + x2 = 0 + 4 Combine like terms: 0 + 4 = 4 2x + x2 = 4 The x term is 2x. Take half its coefficient (1). Square it (1) and add it to both sides. Add '1' to each side of the equation. 2x + 1 + x2 = 4 + 1 Reorder the terms: 1 + 2x + x2 = 4 + 1 Combine like terms: 4 + 1 = 5 1 + 2x + x2 = 5 Factor a perfect square on the left side: (x + 1)(x + 1) = 5 Calculate the square root of the right side: 2.236067978 Break this problem into two subproblems by setting (x + 1) equal to 2.236067978 and -2.236067978. Subproblem 1 x + 1 = 2.236067978 Simplifying x + 1 = 2.236067978 Reorder the terms: 1 + x = 2.236067978 Solving 1 + x = 2.236067978 Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '-1' to each side of the equation. 1 + -1 + x = 2.236067978 + -1 Combine like terms: 1 + -1 = 0 0 + x = 2.236067978 + -1 x = 2.236067978 + -1 Combine like terms: 2.236067978 + -1 = 1.236067978 x = 1.236067978 Simplifying x = 1.236067978 Subproblem 2 x + 1 = -2.236067978 Simplifying x + 1 = -2.236067978 Reorder the terms: 1 + x = -2.236067978 Solving 1 + x = -2.236067978 Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '-1' to each side of the equation. 1 + -1 + x = -2.236067978 + -1 Combine like terms: 1 + -1 = 0 0 + x = -2.236067978 + -1 x = -2.236067978 + -1 Combine like terms: -2.236067978 + -1 = -3.236067978 x = -3.236067978 Simplifying x = -3.236067978 Solution The solution to the problem is based on the solutions from the subproblems. x = {1.236067978, -3.236067978}

OpenStudy (anonymous):

would i subtract four and add it to the other side?

OpenStudy (anonymous):

you would need to subtract it from both sides

OpenStudy (anonymous):

which gives you x^2+2x-4=0

OpenStudy (raden_zaikaria):